# Tackling Cyberattacks through AI-based Reactive Systems: A Holistic Review and Future Vision

### v1

@@ IV-E A Dynamic Decision-Making Approach for Intrusion Response in Industrial Control Systems [96]
A crucial aspect of this decision-making approach is the use of a multilevel Bayesian network model, which offers a profound understanding of how cyber-attacks propagate from the cyber domain to the physical domain within the ICS. This modeling allows for a comprehensive representation of the implications of cyber-attacks and their potential impacts on the entire system. By leveraging this multilevel Bayesian network model, the approach can accurately map detected evidence from real-time IDS, thereby constructing a comprehensive candidate security strategy space based on attack surfaces and functional failures.
@@
In conclusion, the presented security decision-making approach for ICS showcases several noteworthy advantages. Its ability to promptly respond to cyber-attacks and ensure a comprehensive candidate security strategy space makes it a valuable tool for bolstering ICS security. The incorporation of fine-tuned candidate security measures through MOO enables a balanced consideration of security benefit, system benefit, and state benefit, effectively optimizing security performance while preserving the overall system’s functionality. However, a potential limitation lies in its focus on the most dangerous attacks, possibly overlooking less conspicuous threats that may still pose risks and be exploited by attackers. To further enhance its effectiveness, future developments could explore more inclusive and adaptable strategies to address a broader range of potential attack scenarios and fortify ICS environments against ever-evolving cyber threats. To conclude, the absence of artificial intelligence in the proposal is noteworthy. The incorporation of AI models has the potential to effectively mitigate adaptability challenges within the strategy. Furthermore, the use of these models offers the advantages of identifying the most dangerous attacks and improving the decision-making process.
@@ IV-G Exploiting attack–defense trees to find an optimal set of countermeasures [99]
The proposed approach introduces a series of lemmas and corollaries that form an algorithm for extracting possible rational behaviors from the ADTree model in a security environment. These behaviors are defined as strategies, characterized by sets of basic actions performed by both the attacker and defender. This algorithm allows for a systematic exploration of the different ways in which the two actors can interact within the ADTree framework.
@@
Within this paper, a methodology is detailed, based on AGs, to serve the purpose of minimizing risk within a system. Employing ADTree appropriate countermeasures are selected. The first comment about this paper has to do with the lack of use of artificial intelligence. However, ILP has been used together with DAG and ADTrees. It is important to be aware of the fact that the use of ADTrees could provide scalability limitations that could be seen as a constraint in scenarios with many assets  [100]. The automation of the modeling methodology through the proposed tool is a noteworthy aspect. Another advantage of using the OSEAD tool relates to the estimation of the cost of each countermeasure. This provides organisations using the framework with the ability to limit countermeasures according to the budget. To further enhance the proposal, the integration of AI could play a pivotal role in tasks such as automating the creation of ADTree models for different network configurations. Additionally, incorporating dynamic reformulation capabilities to account for changes resulting from attacks or the implementation of countermeasures would enhance the overall adaptability and robustness of the approach. It is therefore considered as an interesting research, which may contain several relevant concepts, yet it might become more autonomous through the use of AI.
@@ IV-H Bayesian Decision Network-Based Security Risk Management Framework [101]
The first phase, risk assessment, focuses on modeling network attacks using AG and calculating the temporal probability of successful vulnerability exploitation. This phase is divided into four steps: modeling network attacks, probability calculation, Bayesian Attack Graph (BAG) construction, and impact calculation. The initial step of the proposed approach necessitates gathering information on existing vulnerabilities, network topology, and host connectivity to generate an AG. To facilitate this process, the authors recommend utilizing tools like MulVAL for AG generation. Moving on to the second phase, probability calculation, security administrators must incorporate the exploitation probability for each vulnerability into its respective nodes using CVSS metrics. Transitioning to the third step, the AG is transformed into a BAG by introducing Conditional Probability Tables (CPTs) to each of its nodes, ensuring that the AG remains an acyclic graph. Finally, the impact of vulnerabilities is computed, considering metrics such as confidentiality requirement, integrity requirement, and availability requirement. These calculations provide crucial insights into the potential consequences of exploiting vulnerabilities within the network’s security framework.
@@
The paper introduces an approach based on a Bayesian network methodology. For the purpose of minimising risk in a system, a probabilistic model is used as algorithm. The framework, utilizing BDNs and BAG, effectively minimize security risks by incorporating vital information about vulnerabilities, risk-reducing countermeasures, and their potential impacts. The proposed framework demonstrates promising ideas in preventive security management. Such as the use of CVSS metrics or the CPTs for the transition to BAG. Nevertheless it has some limitations worth considering. One key limitation is its dependence on security administrators for decision-making during mitigation phase. This is a disadvantage of the model, since the aim of a reactive autonomous system is full task automation and non-reliance on human interaction. In addition, the document focuses mainly on preventive security management, while offering very limited insight into the modelling of reactive countermeasures. This aspect may limit its ability to respond and mitigate incidents in real time.
@@ IV-I Dynamic Countermeasure Knowledge for Intrusion Response Systems [102]
This paper focuses on a countermeasure standard proposal. For this, the authors propose a dynamic countermeasure knowledge model. On a slightly negative note, the document focuses only on outlining the format of countermeasures. The development of a countermeasure selection system is left aside. However, this approach is very interesting, as a common countermeasure format is necessary to unify efforts and develop autonomous reaction systems. One of the advantages of this approach over the countermeasure standard is a dynamic countermeasure knowledge base. By adding dynamism to countermeasures it is easier for organisations to maintain a good number of countermeasures. That is, as new types of malware emerge, or as attacks have evolved, countermeasures will also evolve in the same way. And, since it is a shared database, all systems with such a knowledge base will benefit from updates. The only drawback that can be identified with this type of distributed database is the maintenance. That is, who would be in charge of monitoring the countermeasures that are included, as well as the quality of these countermeasures? It is important to do this because if access is freely open, malicious actors could modify countermeasures to be seen as valid while hiding vulnerabilities in the countermeasures.
@@ IV-K Autonomous mitigation of cyber risks in the Cyber–Physical Systems [107]
The ARC is designed to operate in a hybrid model-free framework, offering the capability to respond to cyber-attacks with or without human intervention, depending on the criticality of the CPS assets at risk. The framework incorporates the quantitative Hierarchical Risk Correlation Tree (HRCT), which effectively models attacker paths and quantifies the financial risk that CPS assets may face from cyber-attacks. By leveraging the HRCT, the ARC gains a comprehensive understanding of the potential threats and vulnerabilities within the CPS.
The power systems serve as a prime application domain for the proposed approach, and the ARC exhibits its prowess by delivering a scalable and autonomous response strategy for the CPS. The risk assessment model based on HRCT empowers the ARC to analyze various security enhancement options, evaluating their effectiveness and prioritizing them based on relative benefits and cost justifications. This dynamic decision-making capability ensures that the ARC can adapt and evolve its responses in real time, keeping the CPS protected against both known and emerging threats.
@@
In conclusion, the paper introduces a novel approach for enhancing security by minimizing the risk in CPS. The methodology used for this task is based on the use of HRCTs. The core of this proposal is an ARC algorithm. Several advantages are noteworthy to mention. To begin with, a strong aspect of this framework has to do with the responsiveness of the framework. It is a very attractive idea how the framework is designed to intervene against cyber-attacks based on the criticality of a particular asset. This can be very beneficial for those scenarios where highly critical machines are attacked, and the reaction time does not allow for human intervention. Another positive aspect is related to the generation of long-term security strategies. This could be seen as a two-mode system, in which knowledge is generated and used statically, and when an attack is identified, it is acted upon autonomously and quickly. However, there is a lack of standardisation or proposals for knowledge generated through an attack. Such knowledge could be relevant to generate an intelligence base to improve reaction systems through improved AI.
@@ IV-L A Bio-Inspired Reaction Against Cyberattacks: AIS-Powered Optimal Countermeasures Selection [108]
The authors build upon their previous work [109], which proposed a standard representation for countermeasures in the field. This representation serves as a foundation for their research. The significance of considering the trade-off between the effectiveness of countermeasures and their potential negative impact and cost is emphasized. The importance of selecting an optimal set of countermeasures that achieves the desired remediation while minimizing any adverse effects is highlighted.
@@ IV-M An Intrusion Response Approach for Elastic Applications Based on Reinforcement Learning [110]
Experimental results further validate the efficacy of transfer learning, showcasing its ability to reduce the training time required for model adaptation when instances are removed. Moreover, the study demonstrates the speed-up achieved by the NN-based IRC, particularly in large-scale systems. This controller produces near-optimal sequences of defense actions by effectively utilizing the stateful model of the system based on the MDP framework.
@@ IV-N GTM: Game Theoretic Methodology for optimal cybersecurity defending strategies and investments [112]
The first module (i.e., the Graphical attack engine) uses MulVAL to generate AGs that model all possible attacking scenarios and paths of an organization. It receives as input a vulnerability assessment report that is the output of a vulnerability assessment tool (e.g., Nessus [116]). The produced graph consists of nodes representing logical propositions, where the source of an attacker’s potential privileges is expressed as a propositional expression based on network configuration parameters.
The Data Pool module serves as a centralized repository of guidelines, laws, and reports about cybersecurity and privacy. Its primary objective is to assist organizations in their defense against cyber-attacks. The module functions as both a database and a platform for ongoing enrichment by Chief Information Security Officers (CISOs). These CISOs actively contribute new input, ensuring that the Data Pool remains up-to-date and relevant to the evolving cybersecurity landscape.
@@
The paper [117] introduces a comprehensive approach to countermeasure selection for enhancing the security of critical infrastructures. By utilizing the Meta Attack Language (MAL) [118] framework, the authors propose a methodology that allows for formalizing and addressing countermeasure selection problems effectively. The proposed algorithm, based on a mathematical model, offers flexibility and adaptability to different scenarios, enabling efficient decision-making processes for improving security.
With MAL, IT infrastructure models can be created and reused, providing a structured representation of devices and their relationships. These models serve as the foundation for generating AGs, where nodes depict attack and defense steps, and edges represent the logical flow of attacks.
@@
An iterative approach is employed to simulate attacks and select appropriate countermeasures. Each iteration involves simulating an attack while assuming that the previously generated solution’s defense steps are implemented. Critical steps in the attack paths are identified, and corresponding countermeasures are examined. When a defense step is found that satisfies all constraints, it and its prerequisites are added to the solution. This process continues iteratively, updating the set of countable steps as new defenses are incorporated. By iteratively refining the solution, the research aims to effectively enhance the security of the system while addressing all necessary constraints.
@@
The authors propose a framework for optimal security strategy selection consisting of three steps: risk assessment, construction of a target attack function benefit-cost of protection, and selection of the optimal security strategy. The risk assessment starts by modeling the cyber-attack using an AG, which is then transformed into a Bayesian AG by calculating the time probability of successful exploitation of each vulnerability. To weigh the benefits and costs, the attack benefits and protection costs are quantified, and an objective function is established. The objective function aims to minimize the attack benefits and protection costs while ensuring that the protection cost remains within the allowable limit.
For the selection of the optimal security strategy, the differential evolution algorithm is employed as a population-based optimization technique. It evolves a population of Np D-dimensional parameter vectors through mutation, crossover, and environmental selection operations. The algorithm starts by generating initial test solutions randomly, aiming to find the optimal security strategy by improving the population of solutions based on evolutionary principles. The selection operation determines which test solutions enter the next-generation population as candidate solutions. The iterative process of mutation, crossover, and selection continues until all experimental solutions are selected or until the accuracy of the current solution meets the requirements or a certain number of iterations is reached.
@@ IV-R Discovering Exfiltration Paths Using Reinforcement Learning with Attack Graphs [123]
The method employs multiple terminal states representing different exit nodes of interest and a single initial node. The agent interacts with the network in an episodic manner to determine the optimal exit node based on expected rewards. The tool provides a comprehensive path analysis for cyber operators to identify the best exit node.
@@ IV-T An Intrusion Response System utilizing Deep Q-Networks and System Partitions [126]
A system contains components of different types. These component types form partitions, representing sets of components belonging to the same type. The system operates on states, actions, and rewards, where the reward function considers execution time and cost.
The software execution commences by creating the system model, which is then decomposed into partitions. For each partition, a DNN is generated, and a corresponding agent is trained. These agents are responsible for providing the local near-optimal next action based on the current partition state. The set of predicted optimal local actions collectively leads to a global optimum.
@@ IV-U Cost-damage analysis of attack trees [128]
AT provide a valuable framework for analyzing and strategizing potential attacks, enabling a systematic approach to identify vulnerabilities and mitigate risks. While damage is the primary concern for system owners, understanding the associated cost of an attack provides valuable insights into the feasibility and likelihood of such attacks.
The paper addresses the optimization problem by proposing solutions for both deterministic and probabilistic scenarios, considering different levels of uncertainty. Techniques such as ILP and bottom-up approaches are employed to develop effective methodologies for determining the maximum achievable damage within the given budget.
Additionally, the paper presents three problem statements: finding the most damaging attack given a cost budget, finding the cheapest attack given a damage threshold, and finding the cost-damage Pareto Front. One notable challenge in cost-damage analysis lies in the fact that existing methods often overlook attacks that do not activate the top node of the attack tree, despite their potential to cause significant damage to intermediate nodes. The paper highlights the importance of considering these intermediate attacks in the analysis.
@@
In conclusion, this paper uses AT methodology and efficient algorithms such as heuristic probabilistic methods for solving cost-damage problems. Although the paper does not specifically focus on countermeasure selection systems, it introduces an intriguing concept with practical implications. The findings can be used for understanding security metrics and decision-making processes in the context of budget-constrained attacks. These methods and AT may be deployed in a reaction system for countermeasure selection based on budget. In addition, another advantage of this research has to do with the generation of different attack possibilities. This presents an interesting point of view as it generates information that can be very valuable for the planning of defensive strategies. This can be especially helpful for organisations that have limited budget. By doing so, these organisations could be able to prioritise their resources more effectively and protect the most probable attack scenarios. However, it is important to underline the limitations of AT. Such as the scalability limitations of ADTrees, AT present the same problems as they are based on trees [130].

### v2

@@ IV-E A Dynamic Decision-Making Approach for Intrusion Response in Industrial Control Systems [98]
A key aspect of this approach is the use of a multilevel Bayesian network model, providing insight into how cyber-attacks spread from the cyber to the physical domain within the ICS. This modeling enables a comprehensive representation of the consequences of cyber-attacks on the entire system. By utilizing this model, the approach accurately maps evidence from real-time IDS to construct a comprehensive security strategy space based on attack surfaces and functional failures.
@@
In conclusion, the presented security decision-making approach for ICS showcases several noteworthy advantages. Its ability to promptly respond to cyber-attacks and ensure a comprehensive candidate security strategy space makes it a valuable tool for bolstering ICS security. The incorporation of fine-tuned candidate security measures through MOO enables a balanced consideration of security benefit, system benefit, and state benefit, effectively optimizing security performance while preserving the overall system’s functionality. However, a potential limitation lies in its focus on the most dangerous attacks, possibly overlooking less conspicuous threats that may still pose risks and be exploited by attackers. To further enhance its effectiveness, future developments could explore more inclusive and adaptable strategies to address a broader range of potential attack scenarios and fortify ICS environments against ever-evolving cyber threats. To conclude, the absence of AI in the proposal is noteworthy. The incorporation of AI models has the potential to effectively mitigate adaptability challenges within the strategy. Furthermore, the use of these models offers the advantages of identifying the most dangerous attacks and improving the decision-making process.
@@ IV-G Exploiting attack–defense trees to find an optimal set of countermeasures [101]
Formalizing ADTrees using DAG enables convenient reasoning about attacker actions, facilitating analysis of multiple attacks and corresponding countermeasures. The framework’s flexibility lies in its generic solution using ILP, allowing customization to specific security optimization criteria.
@@
Within this paper, a methodology is detailed, based on AGs, to serve the purpose of minimizing risk within a system. Employing ADTree appropriate countermeasures are selected. The first comment about this paper has to do with the lack of use of AI. However, ILP has been used together with DAG and ADTrees. It is important to be aware of the fact that the use of ADTrees could provide scalability limitations that could be seen as a constraint in scenarios with many assets  [102]. The automation of the modeling methodology through the proposed tool is a noteworthy aspect. Another advantage of using the OSEAD tool relates to the estimation of the cost of each countermeasure. This provides organizations using the framework with the ability to limit countermeasures according to the budget. To further enhance the proposal, the integration of AI could play a pivotal role in tasks such as automating the creation of ADTree models for different network configurations. Additionally, incorporating dynamic reformulation capabilities to account for changes resulting from attacks or the implementation of countermeasures would enhance the overall adaptability and robustness of the approach. It is therefore considered as an interesting research, which may contain several relevant concepts, yet it might become more autonomous through the use of AI.
@@ IV-H Bayesian Decision Network-Based Security Risk Management Framework [103]
The first phase, risk assessment, focuses on modeling network attacks using AG and calculating the temporal probability of successful vulnerability exploitation. This phase is divided into four steps: modeling network attacks, probability calculation, Bayesian Attack Graph (BAG) construction, and impact calculation. The initial step involves gathering information on existing vulnerabilities, network topology, and host connectivity to generate an AG, possibly utilizing tools like MulVAL. Transitioning to the third step, the AG is transformed into a BAG by introducing Conditional Probability Tables (CPTs) to each of its nodes, ensuring that the AG remains an acyclic graph. . Security administrators then incorporate exploitation probabilities into each vulnerability node using CVSS metrics. Finally, the impact of vulnerabilities is computed, considering metrics such as confidentiality requirement, integrity requirement, and availability requirement.
@@
The paper introduces an approach based on a Bayesian network methodology. For the purpose of minimizing risk in a system, a probabilistic model is used as algorithm. The framework, utilizing BDNs and BAG, effectively minimize security risks by incorporating vital information about vulnerabilities, risk-reducing countermeasures, and their potential impacts. The proposed framework demonstrates promising ideas in preventive security management. Such as the use of CVSS metrics or the CPTs for the transition to BAG. Nevertheless it has some limitations worth considering. One key limitation is its dependence on security administrators for decision-making during mitigation phase. This is a disadvantage of the model, since the aim of a reactive autonomous system is full task automation and non-reliance on human interaction. In addition, the document focuses mainly on preventive security management, while offering very limited insight into the modelling of reactive countermeasures. This aspect may limit its ability to respond and mitigate incidents in real time.
@@ IV-I Dynamic Countermeasure Knowledge for Intrusion Response Systems [104]
This paper focuses on a countermeasure standard proposal. For this, the authors propose a dynamic countermeasure knowledge model. On a slightly negative note, the document focuses only on outlining the format of countermeasures. The development of a countermeasure selection system is left aside. However, this approach is very interesting, as a common countermeasure format is necessary to unify efforts and develop autonomous reaction systems. One of the advantages of this approach over the countermeasure standard is a dynamic countermeasure knowledge base. By adding dynamism to countermeasures it is easier for organizations to maintain a good number of countermeasures. That is, as new types of malware emerge, or as attacks have evolved, countermeasures will also evolve in the same way. And, since it is a shared database, all systems with such a knowledge base will benefit from updates. The only drawback that can be identified with this type of distributed database is the maintenance. That is, who would be in charge of monitoring the countermeasures that are included, as well as the quality of these countermeasures? It is important to do this because if access is freely open, malicious actors could modify countermeasures to be seen as valid while hiding vulnerabilities in the countermeasures.
@@ IV-K Autonomous mitigation of cyber risks in the Cyber–Physical Systems [109]
The ARC is designed to operate in a hybrid model-free framework, responding to cyber-attacks autonomously or with human intervention based on CPS asset criticality. It incorporates a quantitative Hierarchical Risk Correlation Tree (HRCT) to model attacker paths and quantify financial risks to CPS assets.
In power systems, the ARC delivers a scalable and autonomous response strategy, analyzing security enhancement options and prioritizing them based on benefits and cost. This dynamic decision-making capability ensures that the ARC can adapt and evolve its responses in real time.
@@
In conclusion, the paper introduces a novel approach for enhancing security by minimizing the risk in CPS. The methodology used for this task is based on the use of HRCTs. The core of this proposal is an ARC algorithm. Several advantages are noteworthy to mention. To begin with, a strong aspect of this framework has to do with the responsiveness of the framework. It is a very attractive idea how the framework is designed to intervene against cyber-attacks based on the criticality of a particular asset. This can be very beneficial for those scenarios where highly critical machines are attacked, and the reaction time does not allow for human intervention. Another positive aspect is related to the generation of long-term security strategies. This could be seen as a two-mode system, in which knowledge is generated and used statically, and when an attack is identified, it is acted upon autonomously and quickly. However, there is a lack of standardization or proposals for knowledge generated through an attack. Such knowledge could be relevant to generate an intelligence base to improve reaction systems through improved AI.
@@ IV-L A Bio-Inspired Reaction Against Cyberattacks: AIS-Powered Optimal Countermeasures Selection [110]
The authors build upon their previous work [111], which proposed a standard representation for countermeasures in the field. This representation serves as a foundation for their research. The significance of considering the trade-off between the effectiveness of countermeasures and their potential negative impact and cost is emphasized.
@@ IV-N GTM: Game Theoretic Methodology for optimal cybersecurity defending strategies and investments [114]
The first module (i.e., the Graphical attack engine) uses MulVAL to generate AGs that model all possible attacking scenarios and paths of an organization. It receives as input a vulnerability assessment report that is the output of a vulnerability assessment tool (e.g., Nessus [118]).
The Data Pool module serves as a centralized repository of guidelines, laws, and reports about cybersecurity and privacy. Its primary objective is to assist organizations in their defense against cyber-attacks. The module functions as both a database and a platform for ongoing enrichment by Chief Information Security Officers (CISOs).
@@
The paper [119] introduces a comprehensive approach to countermeasure selection for enhancing the security of critical infrastructures. By utilizing the Meta Attack Language (MAL) [120] framework, the authors propose a methodology that allows for formalizing and addressing countermeasure selection problems effectively.
@@
An iterative approach is employed to simulate attacks and select appropriate countermeasures. Each iteration involves simulating an attack while assuming that the previously generated solution’s defense steps are implemented. Critical steps in the attack paths are identified, and corresponding countermeasures are examined. When a defense step is found that satisfies all constraints, it and its prerequisites are added to the solution. This process continues iteratively, updating the set of countable steps as new defenses are incorporated.
@@
The authors propose a framework for optimal security strategy selection consisting of three steps: risk assessment, construction of a target attack function benefit-cost of protection, and selection of the optimal security strategy. The risk assessment starts by modeling the cyber-attack using an AG, which is then transformed into a Bayesian AG by calculating the time probability of successful exploitation of each vulnerability. Benefits and costs are quantified to establish an objective function aiming to minimize attack benefits and protection costs within allowable limits.
For the selection of the optimal security strategy, the differential evolution algorithm is employed as a population-based optimization technique. It iteratively evolves a population of Np D-dimensional parameter vectors through mutation, crossover, and selection operations, aiming to improve solutions based on evolutionary principles. The selection operation determines which test solutions enter the next-generation population as candidate solutions. The iterative process of mutation, crossover, and selection continues until all experimental solutions are selected or until the accuracy of the current solution meets the requirements or a certain number of iterations is reached.
@@ IV-R Discovering Exfiltration Paths Using Reinforcement Learning with Attack Graphs [125]
The method employs multiple terminal states representing different exit nodes of interest and a single initial node. The agent interacts with the network in an episodic manner to determine the optimal exit node based on expected rewards.
@@ IV-T An Intrusion Response System utilizing Deep Q-Networks and System Partitions [128]
The system operates on states, actions, and rewards, where the reward function considers execution time and cost.
The software execution commences by creating the system model, which is then decomposed into partitions. For each partition, a DNN is generated, and a corresponding agent is trained. These agents are responsible for providing the local near-optimal next action based on the current partition state.
@@ IV-U Cost-damage analysis of attack trees [130]
AT provides a valuable framework for analyzing and strategizing potential attacks, enabling a systematic approach to identify vulnerabilities and mitigate risks. While damage is the primary concern for system owners, understanding the associated cost of an attack provides valuable insights into the feasibility and likelihood of such attacks.
Additionally, the paper presents three problem statements: finding the most damaging attack given a cost budget, finding the cheapest attack given a damage threshold, and finding the cost-damage Pareto Front. One notable challenge in cost-damage analysis lies in the fact that existing methods often overlook attacks that do not activate the top node of the attack tree, despite their potential to cause significant damage to intermediate nodes.
@@
In conclusion, this paper uses AT methodology and efficient algorithms such as heuristic probabilistic methods for solving cost-damage problems. Although the paper does not specifically focus on countermeasure selection systems, it introduces an intriguing concept with practical implications. The findings can be used for understanding security metrics and decision-making processes in the context of budget-constrained attacks. These methods and AT may be deployed in a reaction system for countermeasure selection based on budget. In addition, another advantage of this research has to do with the generation of different attack possibilities. This presents an interesting point of view as it generates information that can be very valuable for the planning of defensive strategies. This can be especially helpful for organizations that have limited budget. By doing so, these organizations could be able to prioritize their resources more effectively and protect the most probable attack scenarios. However, it is important to underline the limitations of AT. Such as the scalability limitations of ADTrees, AT present the same problems as they are based on trees [132].